OpenStudy (anonymous):

The first term of a sequence is 1, and the fifth term is 47. If the difference between consecutive terms is constant, what is the second term?

ganeshie8 (ganeshie8):

arithmetic sequence a = 1 t5 = 47

OpenStudy (anonymous):

oh ok... this is looking familiar now...

ganeshie8 (ganeshie8):

in arithmetic sequence, nth term tn = a+(n-1)d t5 = 1 + (5-1) d = 47 1 + 4d = 47 4d = 46 d = ?

OpenStudy (anonymous):

thank you so much. i feel silly knowing that the answer was a lot simpler than i thought haha

ganeshie8 (ganeshie8):

its okay hehe